OpenStudy (anonymous):

integral of (1/(x^2*cos^2(x))+3sin(2x)) dx

OpenStudy (loser66):

\[\int \dfrac{1}{x^2cos^2x}+3sin(2x) dx\] ??

OpenStudy (anonymous):

yes

OpenStudy (loser66):

You don't have elementary result for this.

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ x ^{2}\cos ^{2} }dx+3\times \int\limits \sin(2x)dx\]

OpenStudy (anonymous):

\[\int\limits_{1/(x^2\cos^2(x)) dx}^{?}\]

OpenStudy (anonymous):

this takes too much time to solve &write here u should use rule: \[\int\limits udv=uv-\int\limits vdu\]

OpenStudy (anonymous):

yes i know how to solve the second part of sin , but the first part we should use integration by parts but it doesnt work i dont know how to solve the first part !

OpenStudy (anonymous):

i told u how to solve the first one the rule that i wrote u should use that 1/x^2=u 1/cos^2dx=dv

OpenStudy (anonymous):

the second one is easy the first one take too much time let me write it myself & also u tried yourself

OpenStudy (anonymous):

wait a moment plz

OpenStudy (anonymous):

u should write (1/x^2)*(1+tan^2) do u understand why?

OpenStudy (anonymous):

then with the rule i told u: \[u=\frac{ 1 }{ x ^{2} }\] \[dv=1+\tan ^{2}dx\]

OpenStudy (anonymous):

no i didn t understand

OpenStudy (anonymous):

there is a rule the rule was what i write to u \[\int\limits u \times dv=u \times v -\int\limits v \times du\] its a rule

OpenStudy (anonymous):

from this rule u should find what is U&dv in your question(remember i m solving the first part and this rule is for the firs part)

OpenStudy (anonymous):

(U) is 1/x^2 (dv) is (1+tan^2)dx OK?

OpenStudy (anonymous):

just u should find (du) and (v) then just rewrite in the rule

OpenStudy (anonymous):

\[u=\frac{ 1 }{ x ^{2} }\rightarrow du= -\frac{ 1 }{ x }\] \[dv=(1+\tan ^{2})dx \rightarrow v=\int\limits 1+\tan ^{2}dx=tanx\] got it?

OpenStudy (anonymous):

now just rewrite in the rule(the frame i told u)

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

yourwelcome